#42
A thin, spherical, conducting shell of radius R is
Answer #1

Q = charge on the shell

- V = k Q / R

Q = -VR/k

using conservation of energy

initial Kinetic energy + initial electric potential energy = final electric potential energy

(0.5) m Vo2 + k Qe/r = k Qe / R

(0.5) m Vo2 - VRe/r = - VRe/R

(0.5) m Vo2 = - Ve + VRe/r

Vo2 = (2/m) (Ve) [R/r -1]

Vo = sqrt((2/m) (Ve) [R/r -1])